Problem 8-9 Required rate of return Stock R has a beta of 2.5, Stock S has a beta of 0.5, the required return on an average stock is 10%, and the risk-free rate of return is 5%. By how much does the required return on the riskier stock exceed the required return on the less risky stock? Round your answer to two decimal places.
Answer #1

required return=risk-free rate +Beta*(market rate- risk-free rate)

required return for:

R=5+(10-5)*2.5=17.5%

S=5+(10-5)*0.5=7.5%

Hence difference=17.5%-7.5%

=10%